Twice a first number plus a second number is 42, and the first number minus the second number is-6 . Find the numbers.
Answer by ankor@dixie-net.com(22740)

let the two numbers be a & b
:
Write an equation for each statement, then use elimination
"Twice a first number plus a second number is 42,"
and "the first number minus the second number is-6"
:
2a + b = 42
a - b = -6
--------------adding eliminates b, find a
3a = 36
a = 36/3
a = 12
:
I'll let you find b